# Physics beyond the Standard Model with the DSA-2000

### v1

@@ 4.2 Fifth force
$\mathrm{SNR_{\max,p}}\propto\tilde{\alpha}\sum_{i=1}^{N_{P}}T_{i}^{7/2}\,.$ (4.16)
Constraints on $\tilde{\alpha}$ for PTAs have been derived in Ref. [180] using NANOGrav’s 11-yr dataset, as well as in Ref. [30] by NANOGrav using their 15-yr dataset. Similar to the gravitational interaction case, we derive the upper limits on $\tilde{\alpha}$ by translating the result in Ref. [30] using Eq. (4.16), and the stacked parameters of DSA-2000 and NANOGrav
$\frac{\tilde{\alpha}}{\tilde{\alpha}^{(\mathrm{NG-15})}}=\frac{\sum_{j}^{N_{P}^{\mathrm{NG-15}}}(T^{(\mathrm{NG-15})}_{j})^{7/2}}{\sum_{i}^{N_{P}}T_{i}^{7/2}}\,,$ (4.17)
where $\tilde{\alpha}^{(\mathrm{NG-15})}$ is the upper limit on the effective fifth force strength derived by NANOGrav in their 15-yr dataset assuming $f_{\mathrm{DM}}=1$ [30]. The result is shown in the second panel of Fig. 10. We find that with the combined data of NANOGrav and DSA-2000, an order of magnitude improvement on the constraints on $\tilde{\alpha}$ is expected by 2033, and more than two orders of magnitude improvement by 2053.
@@ 5 Neutrino mass measurements with fast radio bursts
Cosmological constraints on the neutrino masses are derived from their influence on the spatial distribution of structures in the Universe, and its background evolution (see, e.g., [201, 202, 203] for review). Due to their large abundance, neutrinos contribute to the overall matter density; however, their velocity dispersion is very small, enabling them to escape potential wells, washing out structures below a certain scale, called the free-streaming scale (see, e.g., [204]). The sum of the neutrino masses, $\sum m_{\nu}$, has been constrained by the Planck mission [205] and in combination with galaxy surveys such as BOSS [206] and DESI [207]. These cosmological measurements are approaching sensitivities that probe the minimal neutrino mass hierarchies, which, based on neutrino oscillation measurements [208, 209], predict $\sum m_{\nu}=0.058\,\text{eV}$ in the normal hierarchy, and $\sum m_{\nu}=0.11\,\text{eV}$ in the inverted hierarchy. In particular, measurements of the Dark Energy Spectroscopic Instrument (DESI) inferred a $95\%$ upper bound of neutrino masses of $\sum m_{\nu}<0.072\,\text{eV}$ for a $\sum m_{\nu}>0\,\text{eV}$ prior when combined with CMB data, excluding the inverted mass hierarchy, and showing posteriors that are peaked around $\sum m_{\nu}=0\,\text{eV}$ [210].
@@
It has since been argued that the DESI measurement may even point to a cosmological preference for negative neutrino masses [211, 212], driven by the combination of a low inferred value of the fractional matter density from DESI BAO [213], and excess measured lensing power of CMB photons, opposite to what one would expect as signatures of neutrino mass [211, 214].
@@
In light of these developments, upcoming surveys such as Euclid or the Legacy Survey of Space and Time (LSST) will provide valuable independent insights into the neutrino mass puzzle [see e.g. 215]. These upcoming late universe probes of weak gravitational lensing are sensitive to neutrino masses via their impact on both background evolution, and spatial distribution of structure. However, to fully realise their potential, these surveys rely on the utilisation of highly non-linear scales. It is precisely at these scales that visible matter, i.e. baryons, and various astrophysical processes will significantly affect the distribution of structures within the Universe. Phenomena such as supernova feedback, cooling processes, stellar winds, and feedback from active galactic nuclei can displace baryonic matter, which in turn influences the surrounding DM. This interaction alters the statistical properties of the overall matter distribution. These collective influences are referred to as baryonic feedback (BF) [see e.g. 216, 217, 218]. Understanding and accounting for these complex interactions is crucial for accurately interpreting data from cosmological surveys, as they obscure and mimic the weak gravitational lensing signal.
@@
Cosmic shear is the weak gravitational lensing effect induced by the LSS on background galaxies [see 219, 220, for reviews]. The primary observable is the shear, $\gamma$, which can be expressed as the second derivative of the scalar lensing potential, itself the line-of-sight projection of the gravitational potential [^12] More precisely, lensing responds to metric perturbations in both the time-time and space-space components of the metric, i.e. the Bardeen potentials. In General Relativity and in absence of anisotropic stress, however, they are equal.. The statistical properties of the shear field and the convergence, $\kappa$, on the two-point level, are the same. By relating the potential to the density contrast via Poisson’s equation, the convergence at redshift $z$ and angular position $\hat{\boldsymbol{n}}$ can be written as (assuming a spatially flat Universe):
[…]
@@
By averaging over an ensemble of sources distributed over redshift, one obtains the effective convergence by suitably rearranging integrations limits:
[…]
@@
[…]
The splitting of the entire source sample into subsamples is called tomography [221] and is a way to restore information about the redshift evolution of structures. Typically, the information almost saturates after splitting the sample into six tomographic bins [222].
@@
$\mathcal{D}_{\mathrm{tot}}(\hat{\boldsymbol{x}},z)=\mathcal{D}_{\mathrm{host}}(z)+\mathcal{D}_{\mathrm{MW}}(\hat{\boldsymbol{x}})+\mathcal{D}_{\mathrm{LSS}}(z,\hat{\boldsymbol{x}})\,.$ (5.6)
@@
[…]
with the proton mass $m_{p}$, the Hubble constant $H_{0}$, the baryon density parameter today $\Omega_{b}$, the electron fraction $\chi_{e}$ and the fraction of electrons in the inter galactic medium $f_{\mathrm{IGM}}$ which we model following [223] (see also [224] for an updated measurement). Lastly, $\delta_{\mathrm{e}}$ is the electron density contrast.
The mean of Equation 5.7, $\langle\mathcal{D}_{\mathrm{LSS}}\rangle(z)$ is an actual observable and is often referred to as the dispersion measure-redshift relation or the Marcquart relation. Given observations of $N_{\mathrm{FRBs}}$, with redshifts $z_{\alpha}$, positions $\hat{\boldsymbol{n}}_{\alpha}$, and total dispersion measures $\mathcal{D}_{\mathrm{tot},\alpha}$, it contains information about the mean baryon density and the expansion history of the Universe [see e.g. 225, 226, 223, 227, 228, 229]. Measuring $\langle\mathcal{D}_{\mathrm{LSS}}\rangle(z)$ requires the specification of the MW component, for which there are different models available [230, 231, 232]. Furthermore, the host component must be marginalised over for unbiased results [227]. The covariance is given by
@@
assuming that the contributions in Equation 5.6 are uncorrelated. For the MW contribution, we simply assume a constant uncertainty, independent of the FRB with $\sigma_{\mathrm{MW},\alpha}=30\;\mathrm{pc}\,\mathrm{cm}^{-3}$. The host contribution’s uncertainty is assumed to be $\sigma_{\mathrm{host},\alpha}=100/(1+z_{\alpha})\;\mathrm{pc}\,\mathrm{cm}^{-3}$ (as also found in [224]) , also uncorrelated between FRBs. Lastly, the covariance induced by the LSS is described in [3] [^13] An implementation of the covariance is provided on https://github.com/rreischke/frb_covariance..
Following the procedure of cosmic shear, we can also average over ensembles of FRBs distributed as $n^{\mathrm{FRB}}_{i}(z)$, so that the averaged, projected dispersion measure is given by [233, 234]:
[…]
@@
[…]
@@
assuming the Limber approximation [235]. Here $W_{A}$ are probe-specific weights and $P_{AB}$ is the three-dimensional power spectrum of the two probes under consideration. In the case considered here, we need the electron, matter and electron $\times$ matter power spectra. For this, we use the pyhmcode [236, 237] which was calibrated to the BAHAMAS simulations and emulated in [238]. The angular power spectrum gives the variance of the spherical harmonic modes of the project field, $a_{\ell m}$, as a function of inverse angular scale $\ell$. Due to the symmetries of the cosmological model, statistical properties of different $\ell$ do not couple in this estimator and the fluctuations are independent of $m$. Any observed spectrum, $\hat{C}_{AB}(\ell)$ will contain a white noise component due to the discreteness of the traced field:
@@
$N_{A}=\left\{\begin{array}[]{cl}\frac{\sigma^{2}_{\epsilon,i}}{2\bar{n}^{s}_{i}}&\mathrm{if\;}A=\kappa_{i}\\ \frac{\sigma^{2}_{\mathrm{host}}}{\bar{n}^{\mathrm{FRB}}_{i}(1+\bar{z}_{i})}&\mathrm{if\;}A=\mathcal{D}_{i}\,,\end{array}\right.$ (5.13)
@@
Measuring the angular power spectra is limited by cosmic variance and pure shot noise. Cosmic variance is simply the sample variance of cosmology and originates from the fact that we observe only one realisation of the Universe. For example, the estimation of the angular power spectrum at $\ell=1$ will only be possible using three independent modes: $m=-1,0,1$, limiting further shrinking of the error. With this in mind, the Gaussian covariance between $\hat{C}_{AB}(\ell)$ and $\hat{C}_{CD}(\ell^{\prime})$ can be calculated by means of Wick’s theorem:
@@ 5.1 Cosmological inference of the sum of the neutrino masses
We create mock noisy data vectors for the fiducial cosmology specified in Table 5. We then sample the posterior distribution over the parameter and priors outlined in Table 5 using emcee [240, 241]. Our interest here is the sum of the neutrino masses, we therefore only show the marginalised constraints on $\sum m_{\nu}$. For the analysis here, we distinguish four scenarios and display the results in Figure 12:
@@
    Cosmic shear only ($\langle\kappa\kappa\rangle$, blue) assuming a Stage IV cosmic shear survey with optimistic settings as described in [239].
@@
Our findings indicate that FRBs can effectively provide the same enhancement in constraining power as adding CMB measurements to a Stage IV cosmic shear survey. This enhancement is primarily driven by the breaking of degeneracies involving baryonic feedback, as demonstrated in [242], alongside additional constraints on the Hubble constant and baryon density. Alternative methods to measure BF are the Sunyaev-Zel’dovich (SZ) effects [243, 244]. In [237] the thermal SZ effect was used to calibrate the feedback strength and breaking degeneracies. Alternatively, the kinetic SZ effect was employed in [245, 246, 247, 248] and has also the potential to constrain feedback models [249, 250].
The $68\,\%$ upper limit on neutrino mass is improved by approximately a factor of three when incorporating 50,000 FRBs into a cosmic shear analysis. This competitive constraint, entirely independent of the CMB, introduces a high degree of complementarity to the measurements, which is essential for consistent constraints in an era of precision cosmology, especially in light of increasing tensions [see 251, for a review].
@@
$h$ reduced Hubble constant $0.674$ $[0,\infty)$
$\Omega_{b}$ baryon density parameters $0.045$ $[0,\Omega_{\mathrm{m}})$
$\sigma_{8}$ fluctuation amplitude at $8\,\mathrm{Mpc}\,h^{-1}$ $0.83$ $[0,\infty)$
$\sum m_{\nu}\,[\mathrm{eV}]$ sum of neutrino mass $0.05$ $[0,\infty)$
$n_{s}$ spectral index of primordial fluctuations $0.963$ $[-4,4]$
$w_{0}$ constant dark energy equation of state $-1$ $[-1,\infty)$
$w_{a}$ varying dark energy equation of state 00 $[-10,10]$
@@
$\langle\mathcal{D}_{\mathrm{host}}\rangle\,[\mathrm{pc}\,\mathrm{cm}^{-3}]$ mean host contribution $100$ $[0,\infty)$
@@
There are several limitations to consider in this analysis, which we will outline below. We assumed a simplistic one-parameter model for the effect of baryonic feedback. More sophisticated models would include a greater number of parameters (typically around six, as seen in [252, 253, 254]). However, we can anticipate that the degradation of constraints provided by cosmic shear becomes even more pronounced for more complex models of baryonic feedback, making their calibration even more crucial. Moreover, the analysis overlooks systematic effects, but since these are not included for any of the probes, the comparison of the final constraining power remains consistent. This also applies to the general simplicity of the inference using a simple Gaussian likelihood and covariance matrix. For the latter, one might expect that the benefit from FRBs is greater, as the non-Gaussian covariance should be more significant for cosmic shear due to its larger total signal from Stage IV surveys compared to DSA-2000.

### v2

@@ 4.2 Fifth force
$\mathrm{SNR_{\max,p}}\propto\tilde{\alpha}\sum_{i=1}^{N_{P}}T_{i}^{3}\,.$ (4.21)
Constraints on $\tilde{\alpha}$ for PTAs have been derived in Ref. [Gresham:2022biw] using NANOGrav’s 11-yr dataset, as well as in Ref. [NANOGrav:2023hvm] by NANOGrav using their 15-yr dataset. Similar to the gravitational interaction case, we derive the upper limits on $\tilde{\alpha}$ by translating the result in Ref. [NANOGrav:2023hvm] using Eq. (4.21), and the stacked parameters of DSA-2000 and NANOGrav
$\frac{\tilde{\alpha}}{\tilde{\alpha}^{(\mathrm{NG-15})}}=\frac{\sum_{j}^{N_{P}^{\mathrm{NG-15}}}(T^{(\mathrm{NG-15})}_{j})^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3}}}{\sum_{i}^{N_{P}}T_{i}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3}}}\,$, (4.22)
where $\tilde{\alpha}^{(\mathrm{NG-15})}$ is the upper limit on the effective fifth force strength derived by NANOGrav in their 15-yr dataset assuming $f_{\mathrm{DM}}=1$ [NANOGrav:2023hvm]. The result is shown in the second panel of Fig. 10. We find that with the combined data of NANOGrav and DSA-2000, an order of magnitude improvement on the constraints on $\tilde{\alpha}$ is expected by 2033, and two orders of magnitude improvement by 2053.
@@ 5 Neutrino mass measurements with fast radio bursts
Cosmological constraints on the neutrino masses are derived from their influence on the spatial distribution of structures in the Universe, and its background evolution (see, e.g., [lesgourgues_massive_2006, abazajian_neutrinos_2015, gerbino_status_2017] for review). Due to their large abundance, neutrinos contribute to the overall matter density; however, their velocity dispersion is very small, enabling them to escape potential wells, washing out structures below a certain scale, called the free-streaming scale (see, e.g., [marulli_effects_2011]). The sum of the neutrino masses, $\sum m_{\nu}$, has been constrained by the Planck mission [planck_parameters_2018] and in combination with galaxy surveys such as BOSS [ivanov_cosmological_2020] and DESI [2024arXiv240403002D]. These cosmological measurements are approaching sensitivities that probe the minimal neutrino mass hierarchies, which, based on neutrino oscillation measurements [deSalas:2020pgw, Esteban:2020cvm], predict $\sum m_{\nu}=0.058\,\text{eV}$ in the normal hierarchy, and $\sum m_{\nu}=0.11\,\text{eV}$ in the inverted hierarchy. In particular, measurements of the Dark Energy Spectroscopic Instrument (DESI) inferred a $95\%$ upper bound of neutrino masses of $\sum m_{\nu}<0.072\,\text{eV}$ for a $\sum m_{\nu}>0\,\text{eV}$ prior when combined with CMB data, excluding the inverted mass hierarchy, and showing posteriors that are peaked around $\sum m_{\nu}=0\,\text{eV}$ [2024arXiv240403002D].
@@
It has since been argued that the DESI measurement may even point to a cosmological preference for negative neutrino masses [Craig:2024tky, Green:2024xbb], driven by the combination of a low inferred value of the fractional matter density from DESI BAO [Loverde:2024nfi], and excess measured lensing power of CMB photons, opposite to what one would expect as signatures of neutrino mass [Craig:2024tky, Lynch:2025ine].
@@
In light of these developments, upcoming surveys such as Euclid or the Legacy Survey of Space and Time (LSST) will provide valuable independent insights into the neutrino mass puzzle [see e.g. euclid_sensitivity_2024]. These upcoming late universe probes of weak gravitational lensing are sensitive to neutrino masses via their impact on both background evolution, and spatial distribution of structure. However, to fully realise their potential, these surveys rely on the utilisation of highly non-linear scales. It is precisely at these scales that visible matter, i.e. baryons, and various astrophysical processes will significantly affect the distribution of structures within the Universe. Phenomena such as supernova feedback, cooling processes, stellar winds, and feedback from active galactic nuclei can displace baryonic matter, which in turn influences the surrounding DM. This interaction alters the statistical properties of the overall matter distribution. These collective influences are referred to as baryonic feedback (BF) [see e.g. chisari_modelling_2019, vogelsberger_cosmological_2020, schneider_baryonic_2020]. Understanding and accounting for these complex interactions is crucial for accurately interpreting data from cosmological surveys, as they obscure and mimic the weak gravitational lensing signal.
@@
Cosmic shear is the weak gravitational lensing effect induced by the LSS on background galaxies [see bartelmann_weak_2001, kilbinger_cosmology_2015, for reviews]. The primary observable is the shear, $\gamma$, which can be expressed as the second derivative of the scalar lensing potential, itself the line-of-sight projection of the gravitational potential [^16] More precisely, lensing responds to metric perturbations in both the time-time and space-space components of the metric, i.e. the Bardeen potentials. In General Relativity and in absence of anisotropic stress, however, they are equal.. The statistical properties of the shear field and the convergence, $\kappa$, on the two-point level, are the same. By relating the potential to the density contrast via Poisson’s equation, the convergence at redshift $z$ and angular position $\hat{\boldsymbol{n}}$ can be written as (assuming a spatially flat Universe):
$\kappa(z,\hat{\boldsymbol{n}})=\frac{3\Omega_{m}}{2\chi_{H}^{2}}\int_{0}^{\chi(z)}{d}\chi^{\prime}\frac{\chi^{\prime}(\chi(z)-\chi^{\prime})}{\chi(z)}\frac{\delta(\chi^{\prime}\hat{\boldsymbol{n}},\chi^{\prime})}{a(\chi^{\prime})}\,$, (5.1)
@@
By averaging over an ensemble of sources distributed over redshift, one obtains the effective convergence by suitably rearranging integration limits:
$\kappa_{i}(\hat{\boldsymbol{n}})=\int_{0}^{\infty}{d}z\,n^{s}_{i}(z)\kappa(z,\hat{\boldsymbol{n}})=\int_{0}^{\chi_{H}}{{d}\chi}\,W_{\kappa_{i}}(\chi)\delta(\chi\hat{\boldsymbol{n}},\chi)\,$, (5.3)
@@
$W_{\kappa_{i}}(\chi)=\frac{3\Omega_{m}}{2\chi^{2}_{H}}\frac{\chi}{a(\chi)}\int_{\chi}^{\chi_{H}}{d}\chi^{\prime}n^{s}_{i}(\chi^{\prime})\frac{\chi-\chi^{\prime}}{\chi^{\prime}}\,.$ (5.4)
The splitting of the entire source sample into subsamples is called tomography [hu_power_1999] and is a way to restore information about the redshift evolution of structures. Typically, the information almost saturates after splitting the sample into six tomographic bins [sipp_marvin_2021].
@@
[…]
@@
$\mathcal{D}_{\mathrm{LSS}}(\hat{\boldsymbol{n}},z)=\frac{3\Omega_{b}H_{0}}{8\pi Gm_{p}}\chi_{e}\,\int_{0}^{z}{d}z^{\prime}\,\frac{1+z^{\prime}}{E(z^{\prime})}f_{\mathrm{IGM}}(z^{\prime})\big(1+\delta_{e}(\hat{\boldsymbol{n}},z^{\prime})\big)\,$, (5.7)
with the proton mass $m_{p}$, the Hubble constant $H_{0}$, the baryon density parameter today $\Omega_{b}$, the electron fraction $\chi_{e}$ and the fraction of electrons in the inter galactic medium $f_{\mathrm{IGM}}$ which we model following [macquart_census_2022] (see also [2024arXiv240916952C] for an updated measurement). Lastly, $\delta_{\mathrm{e}}$ is the electron density contrast.
The mean of Equation 5.7, $\langle\mathcal{D}_{\mathrm{LSS}}\rangle(z)$ is an actual observable and is often referred to as the dispersion measure-redshift relation or the Marcquart relation. Given observations of $N_{\mathrm{FRBs}}$, with redshifts $z_{\alpha}$, positions $\hat{\boldsymbol{n}}_{\alpha}$, and total dispersion measures $\mathcal{D}_{\mathrm{tot},\alpha}$, it contains information about the mean baryon density and the expansion history of the Universe [see e.g. zhou_fast_2014, walters_future_2018, macquart_census_2022, hagstotz_new_2022, james_measurement_2022, wu_eight_2022]. Measuring $\langle\mathcal{D}_{\mathrm{LSS}}\rangle(z)$ requires the specification of the MW component, for which there are different models available [cordes_new_2001, yao_new_2017, yamasaki_galactic_2020]. Furthermore, the host component must be marginalised over for unbiased results [hagstotz_new_2022]. The covariance is given by
@@
assuming that the contributions in Equation 5.6 are uncorrelated. For the MW contribution, we simply assume a constant uncertainty, independent of the FRB with $\sigma_{\mathrm{MW},\alpha}=30\;\mathrm{pc}\,\mathrm{cm}^{-3}$. The host contribution’s uncertainty is assumed to be $\sigma_{\mathrm{host},\alpha}=100/(1+z_{\alpha})\;\mathrm{pc}\,\mathrm{cm}^{-3}$ (as also found in [2024arXiv240916952C]) , also uncorrelated between FRBs. Lastly, the covariance induced by the LSS is described in [reischke_cosmological_2023] [^17] An implementation of the covariance is provided on https://github.com/rreischke/frb_covariance..
Following the procedure of cosmic shear, we can also average over ensembles of FRBs distributed as $n^{\mathrm{FRB}}_{i}(z)$, so that the averaged, projected dispersion measure is given by [reischke_probing_2021, reischke_consistent_2022]:
$\mathcal{D}_{\mathrm{LSS},i}(\hat{\boldsymbol{n}})=\int_{0}^{\chi_{H}}\!\!{d}\chi\;W_{\mathcal{D}_{i}}(\chi)\delta_{e}\big[\hat{\boldsymbol{x}},z(\chi)\big]\,$, (5.9)
@@
$W_{\mathcal{D}}(\chi)=\frac{3H_{0}^{2}\Omega_{{b}}}{8\pi Gm_{p}}\chi_{\mathrm{e}}[1+z(\chi)]f_{\mathrm{IGM}}(z)\int_{\chi}^{\chi_{H}}{d}\chi^{\prime}\,n^{\mathrm{FRB}}_{i}(\chi^{\prime})\,.$ (5.10)
@@
assuming the Limber approximation [2008PhRvD..78l3506L]. Here $W_{A}$ are probe-specific weights and $P_{AB}$ is the three-dimensional power spectrum of the two probes under consideration. In the case considered here, we need the electron, matter and electron $\times$ matter power spectra. For this, we use the pyhmcode [mead_hydrodynamical_2020, troster_joint_2022] which was calibrated to the BAHAMAS simulations and emulated in [neumann_fast_2024]. The angular power spectrum gives the variance of the spherical harmonic modes of the project field, $a_{\ell m}$, as a function of inverse angular scale $\ell$. Due to the symmetries of the cosmological model, statistical properties of different $\ell$ do not couple in this estimator and the fluctuations are independent of $m$. Any observed spectrum, $\hat{C}_{AB}(\ell)$ will contain a white noise component due to the discreteness of the traced field:
@@
[…]
@@
Measuring the angular power spectra is limited by cosmic variance and pure shot noise. Cosmic variance is simply the sample variance of cosmology and originates from the fact that we observe only one realisation of the Universe. For example, the estimation of the angular power spectrum at $\ell=1$ will only be possible using three independent modes: m=1,0,1m=-1,0,1, limiting further shrinking of the error. With this in mind, the Gaussian covariance between $\hat{C}_{AB}(\ell)$ and $\hat{C}_{CD}(\ell^{\prime})$ can be calculated by means of Wick’s theorem:
@@ 5.1 Cosmological inference of the sum of the neutrino masses
We create mock noisy data vectors for the fiducial cosmology specified in Table 5. We then sample the posterior distribution over the parameter and priors outlined in Table 5 using emcee [goodman_2010, foreman_emcee_2013]. Our interest here is the sum of the neutrino masses, we therefore only show the marginalised constraints on $\sum m_{\nu}$. For the analysis here, we distinguish four scenarios and display the results in Figure 12:
@@
    Cosmic shear only ($\langle\kappa\kappa\rangle$, blue) assuming a Stage IV cosmic shear survey with optimistic settings as described in [blanchard_euclid_2020].
@@
Our findings indicate that FRBs can effectively provide the same enhancement in constraining power as adding CMB measurements to a Stage IV cosmic shear survey. This enhancement is primarily driven by the breaking of degeneracies involving baryonic feedback, as demonstrated in [reischke_calibrating_2023], alongside additional constraints on the Hubble constant and baryon density. Alternative methods to measure BF are the Sunyaev-Zel’dovich (SZ) effects [sunyaev_observations_1972, sunyaev_microwave_1980]. In [troster_joint_2022] the thermal SZ effect was used to calibrate the feedback strength and breaking degeneracies. Alternatively, the kinetic SZ effect was employed in [Hand:2012ui, Planck:2015ywj, DES:2016umt, PhysRevD.93.082002] and has also the potential to constrain feedback models [PhysRevD.103.063513, 2021PhRvD.103f3514A].
The $68\,\%$ upper limit on neutrino mass is improved by approximately a factor of three when incorporating 50,000 FRBs into a cosmic shear analysis. This competitive constraint, entirely independent of the CMB, introduces a high degree of complementarity to the measurements, which is essential for consistent constraints in an era of precision cosmology, especially in light of increasing tensions [see abdalla_cosmology_2022, for a review].
@@
$h$ reduced Hubble constant $0.674$ $0,\infty)$
$\Omega_{b}$ baryon density parameters $0.045$ $0,\Omega_{\mathrm{m}})$
$\sigma_{8}$ fluctuation amplitude at $8\,\mathrm{Mpc}\,h^{-1}$ $0.83$ $0,\infty)$
$\sum m_{\nu}\,[\mathrm{eV}]$ sum of neutrino mass $0.05$ $0,\infty)$
$n_{s}$ spectral index of primordial fluctuations $0.963$ [4,4][-4,4]
$w_{0}$ constant dark energy equation of state 1-1 $1,\infty)$
$w_{a}$ varying dark energy equation of state 0 [10,10][-10,10]
@@
$\langle\mathcal{D}_{\mathrm{host}}\rangle\,[\mathrm{pc}\,\mathrm{cm}^{-3}]$ mean host contribution $100$ $0,\infty)$
@@
There are several limitations to consider in this analysis, which we will outline below. We assumed a simplistic one-parameter model for the effect of baryonic feedback. More sophisticated models would include a greater number of parameters (typically around six, as seen in [schneider_new_2015, arico_bacco_2021, giri_emulation_2021]). However, we can anticipate that the degradation of constraints provided by cosmic shear becomes even more pronounced for more complex models of baryonic feedback, making their calibration even more crucial. Moreover, the analysis overlooks systematic effects, but since these are not included for any of the probes, the comparison of the final constraining power remains consistent. This also applies to the general simplicity of the inference using a simple Gaussian likelihood and covariance matrix. For the latter, one might expect that the benefit from FRBs is greater, as the non-Gaussian covariance should be more significant for cosmic shear due to its larger total signal from Stage IV surveys compared to DSA-2000.